# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

### v1

@@ 5 Applications of Machine Learning in Agriculture
ML techniques have gained significant popularity in the field of agriculture, offering promising solutions to enhance farming practices, optimize crop yield, and minimize losses. These techniques can be effectively utilized at various stages of agriculture, as demonstrated in the following sections.
@@ 5.1 Machine Learning Applications in Pre-Harvesting
Numerous studies have explored the application of ML in pre-harvesting. Radhika et al. [41] compared support vector regression (SVR) and multi-layer perceptron (MLP) for atmospheric temperature prediction, highlighting the superior performance of SVR. Srunitha et al. [42] demonstrated the effectiveness of support vector machines (SVM) in accurately classifying non-sandy soils, suggesting the potential for ML to enhance soil property prediction. Nahvi et al. [43] focused on improving the accuracy of estimating soil temperature using the Extreme Learning Machine (ELM) technique with the Self-adaptive evolutionary (SaE) algorithm. Akshay et al. [44] emphasized the importance of precise soil moisture measurement and management for optimal crop yield, proposing the K-Nearest Neighbors (K-NN) algorithm for data analysis. Uzal et al. [45] found that Convolutional Neural Networks (CNNs) outperformed traditional methods for estimating the number of seeds per soybean pod, demonstrating robust behavior against seasonal changes. Reddy et al. [46] proposed an IoT-based smart irrigation system that predicts water requirements using a decision tree algorithm to enhance irrigation efficiency. AI et al. [47] focused on using CNNs for automated crop disease identification, achieving high accuracy and developing a WeChat applet for disease recognition. Mridha et al. [48] proposed a deep learning-based method for early detection and identification of plant diseases, outperforming traditional models and offering a web-based AI tool. Muhammed et al. [49] presented a user-friendly crop recommendation system (UACR) that leverages AIoT architecture to assist farmers in selecting suitable crops based on location, rain level, and soil type.
@@ 5.2 Machine Learning Applications in Harvesting
Significant advancements have been made in utilizing ML for harvesting purposes. Durmucs-Bunlu et al. [50] discussed the development of a mobile autonomous robot designed for agricultural operations, with the goal of enhancing production efficiency through precision farming, disease diagnosis, and soil analysis. De Baerdemaeker et al. [51] introduced Octinion’s fully autonomous picking robot for strawberry cultivation, addressing labor shortages in the industry. This robot effectively identifies and harvests ripe strawberries, ensuring their quality while neatly arranging them in punnets. Zhang et al. [52] presented an improved deep learning-based method for tomato ripeness classification in a harvesting robot. The study explored various techniques for dataset augmentation to enhance accuracy using convolutional neural networks (CNN). Kestur et al. [53] proposed MangoNet, a deep learning system that utilizes a convolutional neural network and contour-based object detection for mango detection and counting in RGB images, enabling estimation of mango yield. Altaheri et al. [54] described the development, testing, and validation of SWEEPER, an autonomous robot for harvesting sweet pepper fruit in greenhouses. SWEEPER employs an industrial arm and advanced sensing technology to perform its harvesting tasks. Furthermore, Bogue et al. [55] provided insights into recent commercial developments in fruit picking robots, focusing specifically on strawberry and other soft fruit harvesting, while examining their commercial prospects.
@@ 5.3 Machine Learning Applications in Post-Harvesting
In the field of post-harvesting, ML has demonstrated its potential in various applications. Al-Hilali et al. [56] address the challenges faced by Saudi Arabian growers by proposing a visionary solution that harnesses the power of computer vision and RGB images. Their groundbreaking system extracts quality features and achieves precise categorization of dates with an impressive 80% accuracy using a back propagation neural network. Purandare et al. [57] present an end-to-end system aimed at reducing post-harvest losses in agriculture. This system provides real-time data, suggestions for optimal harvesting, and predictive analytics for warehouse management. By employing statistical and probabilistic techniques, it mitigates storage and transportation losses while supporting informed decision-making. Chakravarthy et al. [58] introduce an automated system for classifying mangoes based on their condition using image processing techniques. Leveraging MATLAB and the k-means clustering algorithm, this study extracts features from mango images and evaluates the accuracy of the classification process, ultimately enhancing efficiency in fruit grading and classification. Another notable application is the tomato grading machine vision system described by Ireri et al. [59]. This system utilizes RGB images to achieve accurate and reliable grading by detecting calyx and stalk scars, identifying defected regions, and recognizing various grading categories based on color and texture features. With its high accuracy and potential for inline sorting, this system plays a crucial role in maintaining quality standards during tomato production. Duysak et al. [60] employ the K-nearest neighbor (KNN) algorithm and stepped frequency continuous wave radar (SFCWR) to accurately determine grain quantity in a silo. Through the analysis of back-scattering signals, range profiles, and feature combinations, they achieve an impressive $96.71\%$ accuracy, highlighting the effectiveness of ML in assessing grain quantity with precision. Lutz et al. [61] propose a hierarchical model called Boosted Continuous Non-spatial whole Attribute Extraction (BCNAE) for detecting post-harvest losses in sekai-ichi apples. This model comprises three levels: manual separation, shape fitting, and feature refinement through image segmentation. The BCNAE framework demonstrates effective detection and reduction of post-harvest losses, offering precise solutions in apple production.
5.4 Machine Learning Application in Disease Detection
Researchers have extensively utilized ML and image processing techniques to identify diseases in various types of plants during pre-harvesting, harvesting and post-harvesting. By employing these methods, they aim to accurately detect and diagnose plant diseases, enabling timely intervention and effective management strategies.
An automated approach for leaf disease detection in agriculture is proposed in [62] for tomato plants. The model aims to detect diseases at early stages, enhancing crop production. SVM, Random Forest, and Logistic Regression algorithms are compared, with SVM outperforming the others. The study lacks specific details on dataset size and evaluation metrics. Similarly, a ML and computer vision-based apple disease detection system is proposed in [63]. It achieves $98.63\%$ accuracy by segmenting diseased regions and utilizing features like DWT and color histograms.
For accurate and efficient multi-class plant disease detection, a deep learning object detection model is proposed in [64], surpassing the performance of existing models with a mean average precision of $91.2\%$ and an F1-score of $95.9\%$. Another study [65] introduces a deep convolutional neural network (DCNN) for early diagnosis of apple tree leaf diseases (ATLDs). The proposed DCNN achieves $98.82\%$ accuracy, showing fast convergence, fewer parameters, and high robustness. In [66], an accurate identification approach for common apple leaf diseases is proposed using a deep convolutional neural network. It achieves a $97.62\%$ accuracy with reduced parameters. A lightweight convolutional neural network, RegNet, is introduced in [67] for rapid and accurate identification of apple leaf diseases, surpassing other pre-trained models with a $99.23\%$ overall accuracy. The research in [68] presents a CNN model with $19$ layers for accurate classification of Marsonina Coronaria and Apple Scab diseases in apple leaves, achieving $99.2\%$ accuracy on a dataset of $50,000$ images. Another ML-based approach for automated detection of apple tree diseases is proposed in [69], achieving high accuracy of approximately $95\%$ for disease classification. A deep convolutional neural network model is proposed in [70] for real-time detection of apple leaf diseases, achieving an overall accuracy of $99.31\%$ and surpassing other models. The research in [71] utilizes fine-tuned pre-trained CNN models for accurate apple crop disease detection, with DenseNet201 achieving $98.75\%$ accuracy.
In [72], an automated software solution using image processing techniques is proposed for apple leaf disease detection, based on color and texture features. Finally, an algorithm for accurate segmentation of diseased parts in apple leaf images is proposed in [73], achieving an accuracy of $96.4\%$ with the $k$ nearest neighbor classifier. Table LABEL:tab:long provides a comprehensive overview of the diverse applications of ML in the field of agriculture.
Table 5: Application of ML at various stages of agriculture.
@@ 5.4 Machine Learning Application in Disease Detection
1 Soil Classification 1995 The author is trying to give a quantitative soil classification. Nueral Network The effectiveness of this technique is due to its ability to classify soil quantitatively. Requires a large amount of high-quality training data to achieve accurate results [74]
@@
2 Weather prediction 2009 This paper describes the use of SVM for predicting weather. SVM The finding that SVM consistently outperformed MLP suggests that SVM may be a more effective approach for weather prediction tasks. The selection of appropriate parameters is critical for the successful implementation of SVM in weather predictio. [41]
@@
3 Soil Classification 2016 This paper explains SVM based classification of the soil types. Soil SVM Model is showing less accuracy SVM is a powerful ML technique that can handle large and complex datasets. SVM models are sensitive to the choice of kernel function and its associated parameters, which requires careful tuning to optimize performance. [42]
@@
4 Soil Temperature 2016 Soil in this paper the measurement is for soil temperature (ST) by AI is presented. ELM, SaE-ELM, genetic p rogramming (GP) and ANN Higher generalization performance and prediction accuracy compared to traditional algorithms. SVM-based soil classification shows lower accuracy compared to alternative classification methods. [43]
@@
5 Soyabeen pods 2018 38 types of variety of soyabeen pods has been classified CNN, SVM CNNs offer a more accurate and efficient approach for feature extraction compared to traditional methods. Training and fine-tuning CNNs can be computationally expensive, especially for large datasets. [45]
@@
6 Seed test 2019 This paper proposes a method for classifying plant seedlings using a dataset consisting of around 5,000 images representing 960 unique plants from 12 species at different developmental stages. CNN The proposed solution can assist farmers in optimizing their crop production, which can potentially lead to higher yields and profits. Both CNN and SVM models require careful selection of hyperparameters, which is time-consuming process and require expert knowledge [75]
@@
7 Irrigation plaining 2019 The proposed solution will be developed by establishing a distributed wireless sensor network, wherein each region of the farm would be covered by various sensor modules which will be transmitting data on a common server. IoT, SVR, RF, Wireless sendors Sutaiable and computational efficient approach by Iot and ML Reliability concerns due to the possibility of sensor failures or connectivity issues. [76]
@@
8 Crop predicor and weed detector 2020 The aim of the paper is to provide farmers with a user-friendly and straightforward application that can assist them in selecting the appropriate crops for their land and detecting and eliminating weeds in their crop fields. CNN, WSN, naive bayes Precise management of the farm sector can lead to increased productivity and cost savings for farmers. The model require significant initial investment and ongoing maintenance costs for hardware, software, and data collection [77]
@@
9 Seed test 2020 A mobile application has been developed that quickly detects and classifies seed images with high accuracy using CNN, one of the deep learning techniques. CNN, Inceptionv3, Xception and Inception ResNetV2 (i) The mobile application can detect and classify seed images with high accuracy using CNN, a powerful deep learning technique. (ii) The use of advanced models such as Inceptionv3, Xception, and InceptionResNetV2 can enhance the performance of the application. The system not able to identify all seed species accurately, especially rare or less common ones.. [78]
10 Irrigation plaining 2020 The authur has propsed a irrigation system using ML CNN , DT (i) The use of CNN architecture helps in capturing complex patterns and relationships in the data. (ii) The decision tree algorithm helps in improving the interpretability and explainability of the model The model’s performance may be affected by the quality and accuracy of the data used for training, as well as changes in weather conditions [79]
@@
11 Irrigation planing 2020 This paper presents an advanced technology based smart system to predict the irrigation requirements of a field by sensing of ground parameter like moisture of the soil, temperature- humidity and water level using ML algorithm The paper proposes a smart system that uses ML algorithms to predict the irrigation requirements of a field by sensing ground parameters such as soil moisture, temperature-humidity, and water level. RF (i) The system offers a solution to the over fitting problem faced by existing algorithms like K-mean and SVM. (ii) The system has the capacity to realize a fully automated irrigation scheme. The system may not perform optimally under extreme weather conditions or in areas with highly variable soil types [44]
@@
12 Irrigation planing 2020 This research presents a smart irrigation system that utilizes ML algorithms to predict the water requirements of a crop. DT Better yield prediction. Decision trees are prone to over-fitting the training data [46]
13 Disease detection 2020 CNN is used to automatically identify crop diseases Inception-ResNet-v2 This model for automatic crop disease detection save time and reduce human effort. Shows lower accuracy when compared with other ML model [47]
@@
14 Soil 2020 The study aims to develop a model that can predict soil fertility indices for different villages based on soil reaction, organic carbon, and nutrient content. ELM with different activation functions like sine-squared, Gaussian radial basis etc The approach is cost-effective and efficient, as it eliminates the need for expensive and time-consuming laboratory tests. Not able to capture the complex interactions between different soil parameters and their impact on soil fertility indices [80]
@@
15 Crop recommendation system 2021 The paper presents a yield prediction system for farmers that is both practical and user-friendly. SVM, ANN, RF, multi linear regression These models handles large amount of data and complexity of data The effectiveness of crop selection assistance may depend on the availability and accuracy of data on soil and climate conditions [81]
16 disease detection and croprecommendation system 2021 This paper presents location and season-based crop recommendation system CNN i.e. Sequential and VGG-16. The recommendation feature can also help farmers to diversify their crops and reduce the risk of crop failure. Overfitiing when performs on limited dataset [82]
@@
17 Seed test 2021 This paper shows that wheat seed classification. Neural network The model has High accuracy. Neural networks can be complex to design and train, requiring specialized knowledge and resources. [83]
18 Disease detection 2021 The proposed model is a hybrid system based on Convolutional Autoencoder (CAE) network and CNN for automatic plant disease detection. CNN The hybrid model can potentially improve the accuracy of disease detection compared to traditional methods. The effectiveness of the model depend on the quality and quantity of the training dataset. [84]
19 Disease detection 2021 Computer vision and AI can aid in the early detection of plant diseases, providing a solution to the challenges of continuous human monitoring and reducing the negative impacts of diseases on crops. CNN Saves time of farmer Model is dependent on high quality image data [85]
@@
20 Crop recomendation system 2022 The author has presented an analysis of AI-driven precision farming and agriculture, including related studies, and has proposed a new crop recommendation platform that utilizes cloud-based ML. The platform is designed to aid farmers in making decisions regarding which crops to harvest by considering various known factors. KNN, DT, RF, XGBoost and SVM The platform can potentially reduce waste by providing farmers with more accurate information about when to harvest their crops, resulting in fewer spoiled or unsellable crops The platform require a significant amount of data to be effective, which may be difficult to obtain in some regions or for certain types of crops. [86]
@@
21 Crop recomendation system 2022 The proposed technique combines IoT and AI to develop a smart agricultural system. IoT, layer Perceptron neural Multinetwork The system can optimize the use of resources such as water and fertilizers, reducing waste and environmental impact. The system require regular maintenance and updates to ensure optimal performance, which is time-consuming and expensive. [87]
@@
22 Seed test 2022 This research work aims to explore the application of deep learning techniques for seed classification. MobileNetV2 The choice of using the MobileNetV2 model was based on its efficient memory usage and straightforward architecture. Less seed class taken [88]
23 Disease detection 2022 A model using CNN and VGG is developed to accurately classify disease-affected leaves. . CNN, VGG Classify disease-affected leaves, allowing for early detection and treatmenty Time complexity [89]
24 Disease detection 2022 The study employed a hybrid ML model to detect plant diseases and A web-based AI tool has been developed using this model for plant disease identification. CNN, AlexNet and LetNet Found it to be superior to traditional models. Does not work well in large data [48]
@@
25 Soil Classification 2022 The study aimed to assess the effectiveness of ML modelin classifying potential acid sulfate soils in Jutland, Denmark, and also to apply a model-agnostic interpretation method on the resulting CNN model for improved understanding and insights CNN The use of CNNs for the classification of potential acid sulfate soils is a promising approach for contextual mapping in wetland areas The CNN model’s accuracy and interpretability may be influenced by the quality and quantity of training data used. [90],
@@
26 Crop recomendation system 2023 The proposed crop recommendation system, UACR, allows farmers to input their farm location and automatically receive recommendations for the best crops to grow based on various factors such as rain level and soil type. Iot The UACR system simplifies the crop recommendation process for farmers, as they only need to input their location The absence of edge/fog nodes in some architectures may pose challenges for processing the large amounts of dat [49]
@@
27 Irrigation plaining 2023 The research proposes an IoT-based intelligent control system for smart agriculture that utilizes sensors to collect environmental data and includes an automated irrigation system. KNN, SVM, DT, RF The system can provide real-time monitoring of environmental conditions Large computing costs. [91]
@@
28 Intrusion detection system 2023 The paper introduces a novel system for detecting intrusions in IoT networks utilized in agriculture, using deep learning models and image-based features CNN, PSO for optimization, VGG16, Exception, and Inception model Evaluating the proposed method with different models can help identify potential weaknesses and areas for improvement. Time-consuming [92]
@@
29 Soil Classification 2023 This study demonstrates the application of ML for soil classification using ML XGBoost,KNN imputer Model gives more accurate result Model is limited to data quality [93]
@@
30 Pepper seeds 15 features classified based on width , length, and projected area, seed weight and densityusing ML MLP, BLR, single feature models Model has high accuracy in classification Less accuracy on more features [94]
@@
31 Mobile autonomous robot for agricultural operations 2015 This technology enables accurate analysis and classification of fruit images, including those with overlapping fruits, using semantic segmentation for individual fruits and instance segmentation for overlapping fruits. R-CNN Models saves time of labour High cost of robots [50]
@@
32 Fruit Classification 2018 We propose an efficient framework for fruit classification using deep learning. VGG-16 fine-tuned model, CNN The system can potentially reduce food waste by accurately identifying and classifying fruits, The initial cost of implementing the system high [95]
@@
33 A fully autonomous picking robot 2018 The paper presented AI robot for fruits AI Perception Alternative for costly human picker with no damage Ensure robustness and to identify its shortcomings. [51]

### v2

@@ 3 Applications of Machine Learning in Agriculture
To address RQ-1, the application of ML techniques in agriculture is extensively examined, highlighting their potential to enhance farming practices, optimize crop yields, and reduce losses. The effectiveness of these techniques across various stages of agriculture is demonstrated as follows:
@@ 3.1 Machine Learning Applications in Pre-Harvesting
Numerous studies have explored the application of ML in pre-harvesting. For example, Radhika et al. [79] compared support vector regression (SVR) and multi-layer perceptron (MLP) for atmospheric temperature prediction, highlighting the superior performance of SVR. Srunitha et al. [122] demonstrated the effectiveness of support vector machines (SVM) in accurately classifying non-sandy soils, suggesting the potential for ML to enhance soil property prediction. Nahvi et al. [140] focused on improving the accuracy of estimating soil temperature using the Extreme Learning Machine (ELM) technique with the Self-Adaptive Evolutionary (SaE) algorithm. Akshay et al. [128] emphasized the importance of precise soil moisture measurement and management for optimal crop yield, proposing the K-Nearest Neighbors (K-NN) algorithm for data analysis. Uzal et al. [145] found that CNNs outperformed traditional methods to estimate the number of seeds per soybean pod, demonstrating robust behavior against seasonal changes. Reddy et al. [130] proposed an IoT-based smart irrigation system that predicts water requirements using a decision tree algorithm to improve irrigation efficiency. Ai et al. [134] focused on using CNNs for automated crop disease identification, achieving high precision, and developing a WeChat applet for disease recognition. Mridha et al. [135] proposed a deep learning-based method for early detection and identification of plant diseases, outperforming traditional models and offering a web-based AI tool. Muhammed et al. [136] presented a user-friendly crop recommendation system (UACR) that uses AIoT architecture to assist farmers in selecting suitable crops based on location, rain level and soil type.
@@ 3.2 Machine Learning Applications in Harvesting
Significant advances have been made in the use of ML for harvesting purposes. Durmucs-Bunlu et al. [54] discussed the development of a mobile autonomous robot designed for agricultural operations, with the goal of enhancing production efficiency through precision farming, disease diagnosis, and soil analysis. De Baerdemaeker et al. [62] introduced Octinion’s fully autonomous picking robot for strawberry cultivation, addressing labor shortages in the industry. This robot effectively identifies and harvests ripe strawberries, ensuring their quality and neatly arranging them in punnets. Zhang et al. [64] presented an improved deep learning-based method for classification of tomato ripeness in a harvesting robot. The study explored various techniques for augmentation of data sets to improve accuracy using CNN. Kestur et al. [61] proposed MangoNet, a deep learning system that utilizes a convolutional neural network and contour-based object detection for mango detection and counting in RGB images, enabling the estimation of mango yield. Altaheri et al. [213] described the development, testing and validation of SWEEPER, an autonomous robot that is used to harvest sweet pepper fruits in greenhouses. SWEEPER employs an industrial arm and advanced sensing technology to perform its harvesting tasks. Furthermore, Bogue et al. [53] provided information on recent commercial developments in fruit picking robots, focusing specifically on strawberry and other soft fruit harvesting, while examining their commercial prospects.
@@ 3.3 Machine Learning Applications in Post-Harvesting
In the field of post-harvesting, ML has demonstrated its potential in various applications. Al-Hilali et al. [73] address the challenges faced by Saudi Arabian growers by proposing a visionary solution that harnesses the power of computer vision and RGB images. Their groundbreaking system extracts quality features and achieves precise categorization of dates with an impressive 80% accuracy using a back-propagation neural network. Purandare et al. [74] present an end-to-end system aimed at reducing post-harvest losses in agriculture. This system provides real-time data, suggestions for optimal harvesting, and predictive analytics for warehouse management. Using statistical and probabilistic techniques, it mitigates storage and transportation losses while supporting informed decision-making. Chakravarthy et al. [70] introduce an automated system to classify mangoes according to their condition using image processing techniques. Using MATLAB and the k-means clustering algorithm, this study extracts characteristics from mango images and evaluates the precision of the classification process, ultimately improving efficiency in the classification and classification of fruits. Another notable application is the tomato grading machine vision system described by Ireri et al. [72]. This system utilizes RGB images to achieve accurate and reliable grading by detecting calyx and stalk scars, identifying defected regions, and recognizing various grading categories based on color and texture features. With its high accuracy and potential for inline sorting, this system plays a crucial role in maintaining quality standards during tomato production. Duysak et al. [68] employ the K-nearest neighbor (KNN) algorithm and stepped frequency continuous wave radar (SFCWR) to accurately determine grain quantity in a silo. Through the analysis of back-scattering signals, range profiles, and feature combinations, they achieve an impressive $96.71\%$ accuracy, highlighting the effectiveness of ML in assessing grain quantity with precision. Lutz et al. [67] propose a hierarchical model called Boosted Continuous Non-spatial whole Attribute Extraction (BCNAE) for detecting post-harvest losses in sekai-ichi apples. This model comprises three levels: manual separation, shape fitting, and feature refinement through image segmentation. The BCNAE framework demonstrates effective detection and reduction of post-harvest losses, offering precise solutions in apple production. Table LABEL:tab:long provides a comprehensive overview of the various applications of ML in the field of agriculture.
Table 2: Applications of ML at Different Stages of Agriculture (Pre-harvesting, Harvesting, Post-harvesting).
@@
1 Soil Classification 1995 The author is trying to give a quantitative soil classification. $\bullet$ Nueral Network $\bullet$ The effectiveness of this technique is due to its ability to classify soil quantitatively. $\bullet$ Requires a large amount of high-quality training data to achieve accurate results [123]
@@
2 Weather prediction 2009 This paper describes the use of SVM for predicting weather. $\bullet$ SVM $\bullet$ The finding that SVM consistently outperformed MLP suggests that SVM may be a more effective approach for weather prediction tasks. $\bullet$ The selection of appropriate parameters is critical for the successful implementation of SVM in weather prediction. [79]
@@
3 Soil Classification 2016 This paper explains SVM based classification of the soil types. $\bullet$ SVM $\bullet$ Model is showing less accuracy $\bullet$ SVM is a powerful ML technique that can handle large and complex datasets.
$\bullet$ SVM models are sensitive to the choice of kernel function and its associated parameters, which requires careful tuning to optimize performance.
[122]
4 Soil Temperature 2016 Soil in this paper the measurement is for soil temperature (ST) by AI is presented. $\bullet$ ELM
$\bullet$ SaE-ELM
$\bullet$ genetic p rogramming (GP)
$\bullet$ ANN
@@
$\bullet$ Higher generalization performance and prediction accuracy compared to traditional algorithms. $\bullet$ SVM-based soil classification shows lower accuracy compared to alternative classification methods. [140]
5 Soyabeen pods 2018 38 types of variety of soyabeen pods has been classified. $\bullet$ CNN
$\bullet$ SVM
@@
$\bullet$ CNNs offer a more accurate and efficient approach for feature extraction compared to traditional methods. $\bullet$ Training and fine-tuning CNNs can be computationally expensive, especially for large datasets. [145]
@@
6 Seed test 2019 This paper proposes a method for classifying plant seedlings using a dataset consisting of around 5,000 images representing 960 unique plants from 12 species at different developmental stages. $\bullet$ CNN $\bullet$ The proposed solution can assist farmers in optimizing their crop production, which can potentially lead to higher yields and profits. $\bullet$ Both CNN and SVM models require careful selection of hyperparameters, which is a time-consuming process and requires expert knowledge. [127]
@@
7 Irrigation plaining 2019 The proposed solution will be developed by establishing a distributed wireless sensor network, wherein each region of the farm would be covered by various sensor modules which will be transmitting data on a common server. $\bullet$ IoT
$\bullet$ SVR
$\bullet$ RF
$\bullet$ Wireless sendors
$\bullet$ Sutaiable and computational efficient approach by Iot and ML $\bullet$ Reliability concerns due to the possibility of sensor failures or connectivity issues. [148]
@@
8 Crop prediction and weed detection 2020 The aim of the paper is to provide farmers with a user-friendly and straightforward application that can assist them in selecting the appropriate crops for their land and detecting and eliminating weeds in their crop fields. $\bullet$ CNN
$\bullet$ WSN
$\bullet$ naive bayes
$\bullet$ Precise management of the farm sector can lead to increased productivity and cost savings for farmers. $\bullet$ The model requires significant initial investment and ongoing maintenance costs for hardware, software, and data collection. [96]
@@
9 Seed test 2020 A mobile application has been developed that quickly detects and classifies seed images with high accuracy using CNN, one of the deep learning techniques. $\bullet$ CNN
$\bullet$ Inceptionv3
$\bullet$ Xception
$\bullet$ Inception ResNetV2
$\bullet$ The mobile application can detect and classify seed images with high accuracy using CNN, a powerful deep learning technique.
$\bullet$ The use of advanced models such as Inceptionv3, Xception, and InceptionResNetV2 can enhance the performance of the application.
$\bullet$ The system is not able to identify all seed species accurately, especially rare or less common ones. [150]
10 Irrigation plaining 2020 The authur has propsed a irrigation system using ML. $\bullet$ CNN
$\bullet$ DT
$\bullet$ The use of CNN architecture helps to capture complex patterns and relationships in the data.
$\bullet$ The decision tree algorithm helps in improving the interpretability and explainability of the model.
$\bullet$ The model’s performance may be affected by the quality and accuracy of the data used for training, as well as changes in weather conditions. [49]
@@
11 Irrigation planing 2020 This paper presents an advanced technology based smart system to predict the irrigation requirements of a field by sensing of ground parameter like moisture of the soil, temperature-humidity and water level using ML algorithm. $\bullet$ RF $\bullet$ The system offers a solution to the over-fitting problem faced by existing algorithms like K-mean and SVM.
$\bullet$ The system has the capacity to realize a fully automated irrigation scheme.
$\bullet$ The system may not perform optimally under extreme weather conditions or in areas with highly variable soil types. [128]
@@
12 Irrigation planing 2020 This research presents a smart irrigation system that utilizes ML algorithms to predict the water requirements of a crop. $\bullet$ DT $\bullet$ Better yield prediction $\bullet$ Decision trees are prone to over-fitting the training data. [130]
@@
13 Soil 2020 The study aims to develop a model that can predict soil fertility indices for different villages based on soil reaction, organic carbon, and nutrient content. $\bullet$ ELM with different activation functions like sine-squared, Gaussian radial basis etc. $\bullet$ The approach is cost-effective and efficient, as it eliminates the need for expensive and time-consuming laboratory tests. $\bullet$ Not able to capture the complex interactions between different soil parameters and their impact on soil fertility indices. [143]
14 Crop recommendation system 2021 The paper presents a yield prediction system for farmers that is both practical and user-friendly. $\bullet$ SVM
$\bullet$ ANN
$\bullet$ RF
$\bullet$ multi linear regression
@@
$\bullet$ These models handles large amount of data and complexity of data $\bullet$ The effectiveness of crop selection assistance may depend on the availability and accuracy of data on soil and climate conditions. [95]
@@
15 Seed test 2021 This paper shows that wheat seed classification. $\bullet$ Neural network $\bullet$ The model has High accuracy. $\bullet$ Neural networks can be complex to design and train, requiring specialized knowledge and resources. [125]
16 Crop recomendation system 2022 The author has presented an analysis of AI-driven precision farming and agriculture, including related studies, and has proposed a new crop recommendation platform that utilizes cloud-based ML. $\bullet$ KNN
$\bullet$ DT
$\bullet$ RF
$\bullet$ XGBoost
$\bullet$ SVM
@@
$\bullet$ The platform can potentially reduce waste by providing farmers with more accurate information about when to harvest their crops, resulting in fewer spoiled or unsellable crops $\bullet$ The platform requires a significant amount of data to be effective, which may be difficult to obtain in some regions or for certain types of crops. [139]
17 Crop recomendation system 2022 The proposed technique combines IoT and AI to develop a smart agricultural system. $\bullet$ IoT
$\bullet$ layer Perceptron neural Multinetwork
@@
$\bullet$ The system can optimize the use of resources such as water and fertilizers, reducing waste and environmental impact. $\bullet$ The system requires regular maintenance and updates to ensure optimal performance, which is time-consuming and expensive. [138]
@@
18 Seed test 2022 This research work aims to explore the application of deep learning techniques for seed classification. $\bullet$ MobileNetV2 $\bullet$ The choice of using the MobileNetV2 model was based on its efficient memory usage and straightforward architecture. $\bullet$ Less seed class taken [126]
@@
19 Soil Classification 2022 The study aimed to assess the effectiveness of ML modelin classifying potential acid sulfate soils in Jutland, Denmark, and also to apply a model-agnostic interpretation method on the resulting CNN model for improved understanding and insights. $\bullet$ CNN $\bullet$ The use of CNNs for the classification of potential acid sulfate soils is a promising approach for contextual mapping in wetland areas. $\bullet$ The CNN model’s accuracy and interpretability may be influenced by the quality and quantity of training data used. [121]
@@
20 Crop recomendation system 2023 The proposed crop recommendation system, UACR, allows farmers to input their farm location and automatically receive recommendations for the best crops to grow based on various factors such as rain level and soil type. $\bullet$ Iot $\bullet$ The UACR system simplifies the crop recommendation process for farmers, as they only need to input their location. $\bullet$ The absence of edge/fog nodes in some architectures may pose challenges for processing the large amounts of data. [136]
@@
21 Irrigation plaining 2023 The research proposes an IoT-based intelligent control system for smart agriculture that utilizes sensors to collect environmental data and includes an automated irrigation system. $\bullet$ KNN
$\bullet$ SVM
$\bullet$ DT
$\bullet$ RF
$\bullet$ The system can provide real-time monitoring of environmental conditions $\bullet$ Large computing costs. [129]
@@
22 Intrusion detection system 2023 The paper introduces a novel system for detecting intrusions in IoT networks utilized in agriculture, using deep learning models and image-based features. $\bullet$ CNN
$\bullet$ PSO for optimization
$\bullet$ VGG16
$\bullet$ Exception
$\bullet$ Inception model
$\bullet$ Evaluating the proposed method with different models can help identify potential weaknesses and areas for improvement. $\bullet$ Time-consuming [147]
@@
23 Soil Classification 2023 This study demonstrates the application of ML for soil classification using ML. $\bullet$ XGBoost
$\bullet$ KNN imputer
$\bullet$ Model gives more accurate result $\bullet$ Model is limited to data quality [120]
@@
24 Pepper seeds 15 features classified based on width, length, and projected area, seed weight and density using ML. $\bullet$ MLP
$\bullet$ BLR
$\bullet$ single feature models
$\bullet$ Model has high accuracy in classification $\bullet$ Less accuracy on more features [146]
@@
25 Mobile autonomous robot for agricultural operations 2015 This technology enables accurate analysis and classification of fruit images, including those with overlapping fruits, using semantic segmentation for individual fruits and instance segmentation for overlapping fruits. $\bullet$ R-CNN $\bullet$ Models saves time of labour $\bullet$ High cost of robots [54]
@@
26 Fruit Classification 2018 We propose an efficient framework for fruit classification using deep learning. $\bullet$ VGG-16 fine-tuned model
$\bullet$ CNN
$\bullet$ The system can potentially reduce food waste by accurately identifying and classifying fruits $\bullet$ The initial cost of implementing the system is high [168]
@@
27 A fully autonomous picking robot 2018 The paper presented AI robot for fruits. $\bullet$ AI Perception $\bullet$ Alternative for costly human picker with no damage $\bullet$ Ensure robustness and to identify its shortcomings [62]